The Advanced Photon Source
a U.S. Department of Energy Office of Science User Facility

APS-UES Meeting Minutes - 01/12/2016

Attendees: Bruce Glagola, Tom Barkalow, Dean Wyncott, Wendy VanWingeren, Bruno Fieramosca, and Patti Pedergnana

Guest: Elroy Chang

Bruce Glagola

  • The Storage Ring closes on Wednesday January 20th at 0800.
  • X5 Switchgear Maintenance is on Saturday January 16th from 0600-1800.

Tom Barkalow

  • Tom discussed the transformer fire briefly.
  • There are changes being made to the ventilation in the Storage Ring.  They are going to reduce the amount of air being exhausted so that the temperature will rise and hit a stable point, hopefully to keep beam motion to a minimum.

Elroy Chang

  • Elroy provided the LOTO practical to the group.

Bruno Fieramosca

  • Bruno is in the process of calibrating all of the ODH monitors.
  • Bruno is going to contact Yu Huang about the RMD links not working in the shift log.
  • National Elevator will be here for annual certification of chainfalls and elevators.
  • Bruno is also going to ask if our text in the shift log can be edited.

Nena Moonier

  • Lee Makowski’s brain samples have been shipped off site.

Wendy VanWingeren

  • The pre-run briefing will be on Monday, February 1, 2016 at 1400 in the 438 Conference Room.
  • 35-BM is getting a new Front End and Wendy discussed being RSSE for this work.

Patti Pedergnana

  • Scheck’s work is done at Sector 16.  Julie Cross, Mohan, and Eric Rod will be doing a walk through soon.
  • Sector 13 has a new laser being installed in the 434 A020 laboratory.  They received a laser permit yesterday from the ANL LSO.